Please help me with whats wrong?Two crystalline phases of white phosphorus are known. Both contain P4 molecules, but the molecules are packed together in dif
Answer #1

(a) TRUE

\alpha form of the phosphorus gets spontaneously converted to \beta form below -76. 9oC. As the temperature is lowered, the movement of atoms or the randomness will also be lowered. That is, the entropy is decreasing for the process. Therefore the sign of \Delta S will be negative for the process.

(b) FALSE

Given that the above reaction is spontaneous, therefore \Delta G will be negative.

(c) TRUE

\alpha form has a higher entropy than \beta form . As mentioned above, disorder decreases as temperature decreases. Hence \alpha form has a less ordered crystalline structure.

(d) FALSE

\DeltaH for the process is negative.

(e) FALSE

At -76.9oC the system is in equilibrium, therefore \Delta G will be zero.

(f) TRUE

At  -76.9oC the two phases coexist and the system is in equilibrium.